OpenStudy (anonymous):

Solve. x2 + 15x = –56

OpenStudy (anonymous):

x = –8 and x = 7 x = –8 and x = –7 x = 8 and x = 7 x = 8 and x = –7

OpenStudy (anonymous):

Quadratic formula.

OpenStudy (anonymous):

Trial and error checking.

OpenStudy (anonymous):

don't need it, just factor.

OpenStudy (anonymous):

I prefer, \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] Cerezas, you sure this can factor?

OpenStudy (anonymous):

Actually, nevermind. Yes, it can.

OpenStudy (anonymous):

x=-8 and x=-7

OpenStudy (anonymous):

\[x ^{2}+15x+56 =0\] so (x+7)(x+8) = 0 so x = -7, -8

OpenStudy (anonymous):

x^2 + 15x + 56 = 0 Find factors of 56 that add up to 15. (x + 8)(x + 7) = 0 x = -7, -8

OpenStudy (anonymous):

I don't mean to be rude.. but we're supposed to help, not give out answers. "Give Help, Not Answers I will encourage and guide those needing help, and not just give them an answer" http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

That is helpful I'm figuring it out stilll...

OpenStudy (anonymous):

@Limitless it's factoring. there's only so much you can do.

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

@cerezas, I agree. But you could have explained the method and let Katie try it for herself. Or you could've recommended a method. There were several different ways. I think you should have at least let her try and see where her mistakes were so that you could have corrected them.

OpenStudy (anonymous):

@Limitless i did. factoring. it's the first thing to try.

OpenStudy (anonymous):

Yes, but you gave the entire answer... At no point did Katie have to work for herself. (Of course, there is the possibility she had tried this previously. If so, she should've included her work.)

OpenStudy (anonymous):

...or, she can do it herself, using my work as an example. can we take this off her question? we're giving everyone a ton of notifications.

OpenStudy (anonymous):

I just wanted people to know what the rules were. That is all. I personally disagree with people who post just answers. But everyone is entitled to their opinion and I respect that. Thank you for respecting mine.

OpenStudy (anonymous):

sure :)